If the strength of a beam is inversely proportional to the length , and a beam 5 m long can support at most 100 kg , the number of kg an 8 m long beam can support is?
Found 2 solutions by eperette, Mathtut:
Answer by eperette(173)

Concept you should know:
(1)Inverse Proportion y=k/x
where y, x are variables
k is constant of proportionality
***********
S = k/ L
100 = k/5
500 = k
Now given L=8, find S
S = k/L
S = 500/8
S = 62.5 kg

Answer by Mathtut(3670)

lets call the strength of the beam s and the length L
:
sL=k
:
100(5)=500
:
so our constant of variation is 500
:
s(8)=500
:
s=62.5kg